# Frostman random variables, entropy inequalities, and applications

### v3

@@ Abstract
We introduce Frostman conditions for bivariate random variables and study discretized entropy sum–product phenomena in both independent and dependent regimes. We prove that for any non-degenerate rational quadratic form $\phi(x,y)$ and parameter $0<s<1$, there exists a positive constant $\epsilon=\epsilon(\phi,s)$ such that
@@
for sufficiently large $n$, where the precise conditions on $(X,Y)$ depend on the Frostman level. The proof introduces a novel multi‑step entropy framework, combining the submodularity formula, the discretized entropy Balog-Szemerédi-Gowers theorem, and state-of-the-art results on the Falconer distance problem, to reduce general forms to a diagonal core case.
As an application, we derive a result on a discretized sum-product type problem. In particular, for a $\delta$-separated set $A\subset[0,1]$ of cardinality $\delta^{-s}$, satisfying some non-concentration conditions, and a dense subset $G\subset A\times A$, there exists $\epsilon=\epsilon(s,\phi)>0$ such that
@@
for all $\delta$ small enough. Here by $\mathcal{E}_{\delta}(A)$ we mean the $\delta$-covering number of $A$, $\{A+_{G}A:=\{x+y\colon(x,y)\in G\}$, and $\phi_{G}(A):=\{\phi(x,y)\colon(x,y)\in G\}$.
@@ 1 Introduction
For each $n\in\mathbb{N}$, let $D_{n}\colon\mathbb{R}\to\mathbb{R}$ be the function defined by $D_{n}(x)=\frac{k}{2^{n}}$ if $k$ is the unique integer such that $\frac{k}{2^{n}}\leq x<\frac{k+1}{2^{n}}$. For any random variable $X$, $D_{n}(X)$ defines a discrete random variable that takes values in the set of dyadic rationals. We write $H_{n}(X):=H(D_{n}(X))$, where the entropy of a discrete random variable is defined by
@@
In this paper, we establish a general entropy inequality that both sharpens our theoretical understanding and points toward a range of applications. Our study is motivated by recent applications of entropy methods to discretized sum-product estimates [12] due to Máthé and O’Regan, and to the recent solution of the polynomial Freiman-Ruzsa conjecture [2, 3] due to Gowers, Green, Manners, and Tao.
@@
$\max\left\{H_{n}(X+Y),\,H_{n}(X\cdot Y)\right\}\geq n(s+\epsilon).$ (1)
To state our main theorems, we first introduce the following definitions for bivariate random variables $(X,Y)$.
@@
(iii) We say a random variable $X$ is $(s,C)$-Frostman given $Y$ if
@@
for any ball $B(x,r)$ and Borel set $E$. Moreover, we say that the random variable $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, if $X$ is $(s_{1},C_{1})$-Frostman given $Y$, and $Y$ is $(s_{2},C_{2})$-Frostman given $X$.
These definitions are not equivalent. At a common exponent level $(s_{1},s_{2})$, they form a hierarchy:
@@
Our results demonstrate how the Frostman hierarchy affects entropy bounds. We call a quadratic form diagonal if $\phi(x,y)=\alpha x^{2}+\beta y^{2}$, and the main results read as follows.
Theorem 1.2 (Jointly Frostman, diagonal forms).
Let $\alpha,\beta\in{\mathbb{R}}^{\ast}$, $0<s<1$, $C\geq 1$ and $c>0$. There exist a positive constant $\epsilon=\epsilon(s)$ and an integer $N=N(\alpha,\beta,s,C,c)$ such that the following statement holds. For any bivariate real random variable $(X,Y)$ that is jointly $(s_{1},s_{2},C)$-Frostman and has range contained in $c,c]^{2}$, where $0\leq s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$, we have
$\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}(\alpha X^{2}+\beta Y^{2})\right\}\geq n(s+\epsilon)$
@@
Let $\phi(x,y)$ be a non-degenerate rational quadratic form, $0<s<1$ and $C_{1},C_{2},c>0$. There exist positive constants $\epsilon=\epsilon(s)$, $\epsilon^{\prime}=\epsilon^{\prime}(\phi,s)$ and an integer $N=N(\phi,s,C_{1},C_{2},c)$ such that the following statement holds. Let $(X,Y)$ be a bivariate real random variable whose range is contained in $c,c]^{2}$. Suppose that $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, with $0\leq s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$. Then, for all $n>N$,
@@
Here, a constant depends on $\phi$ means that it depends on the coefficients $a_{1}$, $a_{2}$ and $a_{3}$.
@@
Let $\phi(x,y)$ be a non-degenerate rational quadratic form, $0<s<1$ and $C_{1},C_{2},c>0$. There exist positive constants $\epsilon=\epsilon(s)$, $\epsilon^{\prime}=\epsilon^{\prime}(\phi,s)$ and an integer $N=N(\phi,s,C_{1},C_{2},c)$ such that the following statement holds. Let $X$ and $Y$ be independent real random variables whose ranges are contained in [c,c][-c,c]. Suppose that $(X,Y)$ is marginally $(s_{1},C_{1};S_{2},C_{2})$-Frostman, with $0\leq s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$. Then, for all $n>N$,
@@
    The positive $\epsilon$ might not exist if one of the conditions of the theorem is not satisfied, including when $s=0$ or 11 (see Example 7.1), the Frostman constant $C$ grows with $n$ (see Example 7.4), or the polynomial is degenerate (see Example 7.3).
@@
    The value of $\epsilon$ is not explicit, and an inspection of the proofs suggests it would be small.
  7. 7.
    Theorem 1.2 can be extended to a more general form of
    $\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}(f(X)+g(Y))\right\},$
    where $f$ and $g$ are polynomials with $\deg(f)=\deg(g)\geq 2$. Establishing such a result would require extending most of the arguments in this paper, and we plan to provide further updates in a subsequent version.
We pause here to add several further remarks. Our proof of Theorems 1.21.4 presents a connection between $\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}$ and the entropy of the “distance” function, i.e. $H_{n}\big{(}(X_{1}-Y_{1})^{2}+(X_{2}-Y_{2})^{2}\big{)}$, where $(X_{1},Y_{1})$, $(X_{2},Y_{2})$ are i.i.d. copies of $(X,Y)$. When $a_{1}=a_{3}=0$ and $a_{2}=1$, i.e. $\phi(x,y)=xy$, Máthé and O’Regan [12] established a different type of connection by replacing the entropy of the “distance” function with that of the “slope” function. By combining this with a generalization of Marstrand’s theorem due to Orponen, Shmerkin, and Wang in [21], they were able to prove Theorem 1.4 for the full range of $s$ for this specific bilinear form function. We do not see how their method could be adapted to the form of $\phi$ with $a_{1},a_{3}\neq 0$. Moreover, in terms of applications, i.e. sum-product type problems over $\mathbb{R}$ or $\mathbb{F}_{p}$, the two functions $xy$ and $a_{1}x^{2}+a_{2}xy+a_{3}y^{2}$ appear to exhibit distinct behaviors. It would be interesting to explore more connections between $\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}$ and the entropy of other functions.
In the following, we discuss some applications of the entropy theorems.
For $A\subset\mathbb{R}$ and $\delta>0$, by $\mathcal{E}_{\delta}(A)$ we mean the $\delta$-covering number of $A$.
@@
$\alpha\cdot A^{\wedge 2}+_{G}\beta\cdot A^{\wedge 2}=\{\alpha x^{2}+\beta y^{2}:\,(x,y)\in G\}.$
@@
Let $\alpha,\beta\in{\mathbb{R}}^{\ast}$, $0<s<1$ and $C\geq 1$. Then there exist positive constants $\epsilon=\epsilon(s)$ and $\delta_{0}=\delta_{0}(\alpha,\beta,s,C)$ such that the following holds for all $0<\delta\leq\delta_{0}$. Let $A\subset{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}[0,1]}$ be any $\delta$-separated set of cardinality $\lceil\delta^{-s}\rceil$. Suppose that for all intervals $I$ of length $l$ at least $\delta$, we have the non-concentration estimate
@@
$\mathcal{E}_{\delta}(A+_{G}A)+\mathcal{E}_{\delta}(\alpha\cdot A^{\wedge 2}+_{G}\beta\cdot A^{\wedge 2})\gg_{\alpha,\beta,s}C^{-O_{s}(1)}\delta^{-\epsilon}(\#A).$
Thanks to Theorem 1.3, extend this result to all non-degenerate rational quadratic forms is possible, but requiring minimum degree constraints.
@@
Let $\phi(x,y)$ be a non-degenerate rational quadratic form, $0<s<1$ and $C\geq 1$. Then there exist positive constants $\epsilon=\epsilon(\phi,s)$ and $\delta_{0}=\delta_{0}(\alpha,\beta,s,C)$ such that the following holds for all $0<\delta\leq\delta_{0}$. Let $A\subset{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}[0,1]}$ be any $\delta$-separated set of cardinality $\lceil\delta^{-s}\rceil$. Suppose that for all intervals $I$ of length $l$ at least $\delta$, we have the non-concentration estimate
@@
$\{y:\,(x,y)\in G\}\geq\delta^{\epsilon}(\#A)\quad\text{for any}\quad x\in A,$
$\{x:\,(x,y)\in G\}\geq\delta^{\epsilon}(\#A)\quad\text{for any}\quad y\in A,$
@@
$\mathcal{E}_{\delta}(A+_{G}A)+\mathcal{E}_{\delta}(\phi_{G}(A,A))\gg_{\phi,s}C^{-O_{s}(1)}\delta^{-\epsilon}(\#A).$
In contrast to Theorem 1.6, the next result, an application of Theorem 1.4, holds for every scale $\delta>0$ and is stable under taking subsets: it applies to any $B\subset A$ with $B\geq\delta^{\varepsilon}\#A$.
@@
Let $\phi(x,y)$ be a non-degenerate rational quadratic form, $0<s<1$ and $C\geq 1$. Then there exist positive constants $\epsilon=\epsilon(\phi,s)$ such that the following holds for all $\delta>0$. Let $A\subset{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}[0,1]}$ be any $\delta$-separated set of cardinality $\lceil\delta^{-s}\rceil$. Suppose that for all intervals $I$ of length $l$ at least $\delta$, we have the non-concentration estimate
@@
$\mathcal{E}_{\delta}(B+B)+\mathcal{E}_{\delta}(\phi(B,B))\gg_{\phi,s}C^{-O_{s}(1)}\delta^{-\epsilon}(\#A).$
@@
We now sketch the main ideas of proofs in this paper.
Our base case is corresponding to the diagonal forms $\phi(X,Y)=\alpha X^{2}+\beta Y^{2}$ (Lemma 4.1), with independent $(s,C)$-Frostman random variables $X,Y$. Roughly speaking, by applying the submodularity formula (Lemma 2.18) on certain algebraic relations between the sum function $X+Y$ and the distance function $(X-X^{\prime})^{2}+(Y-Y^{\prime})^{2}$, we are able to reduce the problem to bounding the entropy of the distance function from below. In particular, let $W$ be an $(2s,C)$-Frostman random variable with $range(W)\subseteq B(0,R)\subset\mathbb{R}^{2}$, where $0<s<1$, the constant $C\geq 1$ is bounded and does not depend on $n$, and $R>0$. Let $W^{\prime}$ be a random variable which is i.i.d. with $W$. Then there exists a constant $\epsilon_{0}=\epsilon_{0}(s)>0$ such that
$H_{n}(|W-W^{\prime}|^{2})\geq n\left(s+\epsilon_{0}\right)$ (3)
for all $n\geq N_{0}$, where $N_{0}=N_{0}(s,C,R)$. Note that if the Frostman constant $C$ depends on $n$, then the inequality (3) might not be true (see Example 7.5). To prove (3), we make use of results on the Falconer distance problem. While classical estimates taken from [13, Section 4.2] are sufficient in the case $s>1/2$, we invoke the recent work of Shmerkin and Wang [24] for $s\leq 1/2$.
[…]
Then, given that $D_{n}(X^{\prime},Y^{\prime})=(x,y)$ with $x,y$ being dyadic rationals satisfying $P\big{(}D_{n}(X^{\prime},Y^{\prime})=(x,y)\big{)}>0$, in the proof we show that
$\begin{split}\big{(}X^{\prime\prime}|D_{n}(X^{\prime},Y^{\prime})=(x,y)\big{)}\text{ is $\left(s,O\left(C^{2}\cdot 2^{-ns}/P_{E}(D_{n}(Y^{\prime})=y)\right)\right)$-Frostman},\\ \big{(}Y^{\prime\prime}|D_{n}(X^{\prime},Y^{\prime})=(x,y)\big{)}\text{ is $\left(s,O\left(C^{2}\cdot 2^{-ns}/P_{E}(D_{n}(X^{\prime})=x)\right)\right)$-Frostman}.\end{split}$ (4)
The problem now is reduced to the case of random variables $\big{(}X^{\prime\prime}|D_{n}(X^{\prime},Y^{\prime})=(x,y)\big{)}$ and $\big{(}Y^{\prime\prime}|D_{n}(X^{\prime},Y^{\prime})=(x,y)\big{)}$, and the corresponding entropy
$\max\left\{H_{n}\left(X^{\prime\prime}+Y^{\prime\prime}|D_{n}(X^{\prime},Y^{\prime})=(x,y)\right),\,H_{n}\left(\alpha X^{\prime\prime 2}+\beta Y^{\prime\prime 2}|D_{n}(X^{\prime},Y^{\prime})=(x,y)\right)\right\}.$
It is clear that we have many pairs $(x,y)$, and for each pair, the corresponding Frostman constants in (4) are not uniformly bounded. Besides technical issues including Lipschitz conditions and partitions of measure supports, the main difficulty arises because the Frostman constants may grow with $n$. Note that the idea of scaling the Frostman constants to $1$ does not help, since it would change the supports of variables from bounded to unbounded.
We address this challenge by explicitly computing the entropy of the distance function. Instead of (3), we have the following for both small and large dimensions:
Large dimension: Let $W$ be an $(2s,C)$-Frostman random variable with $range(W)\subseteq B(0,R)\subset\mathbb{R}^{2}$, where $1/2<s<1$, $C\geq 1$ and $R>0$. Let $W^{\prime}$ be a random variable which is i.i.d with $W$. Then for any number $t$ satisfying $t\leq 1$ and s<t<2s1/2s<t<2s-1/2, we have that for all $n\in\mathbb{N},$
$H_{n}(|W-W^{\prime}|)\geq nt-\log C+O(|\log R|)+O(|\log(2s-t-1/2)|)+O_{t}(1).$ (5)
Small dimension: Let $X^{\prime\prime}$, $Y^{\prime\prime}$ be independent random variables with ranges contained in [c,c][-c,c], with $c>0$. Suppose that $X^{\prime\prime}$ and $Y^{\prime\prime}$ are $(s,C_{1})$-Frostman and $(s,C_{2})$-Frostman, respectively, where $0<s\leq 1/2$ and $C_{1},C_{2}\geq 1$. Let $(X^{\prime\prime\prime},Y^{\prime\prime\prime})$ be i.i.d. with $(X^{\prime\prime},Y^{\prime\prime})$. Then for any $t$ with $s<t<\frac{s+\varphi(2s)}{2}$, for all $n\in\mathbb{N}$, we have
$H_{n}\big{(}|(X^{\prime\prime},Y^{\prime\prime})-(X^{\prime\prime\prime},Y^{\prime\prime\prime})|\big{)}\geq nt+O_{s}(\log n)+O_{s,t}(\log C_{1}C_{2})+O_{s,c}(1),$ (6)
where
$\varphi(u)=\frac{u}{2}+\frac{\sqrt{4+u^{2}}-2}{2},\quad(0<u\leq 1).$ (7)
By (5) and (6), the error terms involving the Frostman constants in (4) are controlled up to logarithmic level. Then they can be further converted to certain entropy-terms that are feasible to handle. The proof of (5) is essentially the same as that of (3). To prove (6), we adapt the Shmerkin–Wang approach with appropriate modifications, while clarifying, simplifying, and quantifying key steps. Note that our inspection is not verbatim: for instance, in the thin‐tube step we must work with sets having a Cartesian‐product structure. This is why the estimate (6) holds for $W=X^{\prime\prime}\times Y^{\prime\prime}$ rather than in the general form of (5).
Plugging these inequalities into our method, the theorem follows.
Our final comment is that entropy methods have also been used in a range of problems in Additive Combinatorics and Geometric Measure Theory, for example, see [4, 7, 10, 16, 19, 24, 25, 26] and references therein.

### v4

@@ Abstract
We introduce Frostman conditions for bivariate random variables and study discretized entropy sum-product phenomena in both independent and dependent settings. Fix $0<s<1$, and let $(X,Y)$ be a bivariate real random variable with bounded support, whose distribution satisfies a Frostman condition of dimension $s$. Let $\phi(x,y)$ be a polynomial obtained from a diagonal polynomial $\rho_{1}(x)+\rho_{2}(y)\in\mathbb{R}[x,y]$ of degree $d\geq 2$ by applying an invertible rational linear change of variables in $(x,y)$. We show that there exists $\epsilon=\epsilon(\phi,s)>0$ such that
@@
for all sufficiently large $n$, where the precise assumptions on $(X,Y)$ depend on the Frostman level. The proof introduces a novel multi‑step entropy framework, combining the submodularity formula, the discretized entropy Balog-Szemerédi-Gowers theorem, and state-of-the-art results on the Falconer distance problem, to reduce general forms to a diagonal core case.
As an application, we obtain discretized sum-product type estimates. In particular, for a $\delta$-separated set $A\subseteq[0,1]$ of cardinality $\delta^{-s}$, satisfying certain non-concentration conditions, and a dense subset $G\subseteq A\times A$, there exists $\epsilon=\epsilon(s,\phi)>0$ such that
@@
for all $\delta$ small enough. Here $\mathcal{E}_{\delta}(A)$ denotes the $\delta$-covering number of $A$, $A+_{G}A:=\{x+y\colon(x,y)\in G\}$, and $\phi_{G}(A,A):=\{\phi(x,y)\colon(x,y)\in G\}$.
@@ 1 Introduction
For each $n\in\mathbb{N}$, let $D_{n}\colon\mathbb{R}\to\mathbb{R}$ be the function defined by $D_{n}(x)=\frac{k}{2^{n}}$ where $k$ is the unique integer satisfying $\frac{k}{2^{n}}\leq x<\frac{k+1}{2^{n}}$. For any random variable $X$, $D_{n}(X)$ defines a discrete random variable that takes values in the set of dyadic rationals. We write $H_{n}(X):=H(D_{n}(X))$, where the entropy of a discrete random variable is defined by
@@
In this paper, we establish a general entropy inequality that both sharpens our theoretical understanding and points toward a range of applications. Our study is motivated by recent applications of entropy methods to discretized sum-product phenomena [12] due to Máthé and O’Regan, and to the recent solution of the polynomial Freiman-Ruzsa conjecture [2, 3] due to Gowers, Green, Manners and Tao.
@@
$\max\left\{H_{n}(X+Y),\,H_{n}(X\cdot Y)\right\}\geq n(s+\epsilon).$
To state the main theorems, we first introduce the following definitions for bivariate random variables $(X,Y)$.
@@
(iii) We say a random variable $X$ is conditionally $(s,C)$-Frostman given $Y$ if
@@
for any ball $B(x,r)$ and Borel set $E$. Moreover, we say that the random variable $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, if $X$ is conditionally $(s_{1},C_{1})$-Frostman given $Y$, and $Y$ is conditionally $(s_{2},C_{2})$-Frostman given $X$.
These definitions are not equivalent. At a common exponent level $(s_{1},s_{2})$, these conditions form a hierarchy:
@@
The results in this paper demonstrate how the Frostman hierarchy affects entropy bounds.
Theorem 1.2 (Jointly Frostman, diagonal polynomials).
Let $d\geq 2$ and $\rho_{1}$, $\rho_{2}$ be real polynomials of degree $d$. Let $0<s<1$ and $C,c>0$. There exist a positive constant $\epsilon=\epsilon(d,s)$ and an integer $N=N(\rho_{1},\rho_{2},s,C,c)$ such that the following statement holds. For any bivariate real random variable $(X,Y)$ that is jointly $(s_{1},s_{2},C)$-Frostman and has range contained in $c,c]^{2}$, where $1\ll_{s}s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$, we have
$\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}\big(\rho_{1}(X)+\rho_{2}(Y)\big)\right\}\geq n(s+\epsilon)$
@@
In the theorem above, a constant depends on a polynomial $\rho$, we means that it depends on the degree and coefficients of $\rho$. The notation $1\ll_{s}s_{1},s_{2}$ means that $s_{1}$ and $s_{2}$ are bounded below by some positive constant depending only on $s$, but not on $\rho_{1},\rho_{2},C,c$ or $n$.
Next, consider polynomials $\rho_{1}$ and $\rho_{2}$ of degree $d\geq 2$, together with an invertible matrix
$\Xi=\begin{bmatrix}\xi_{11}&\xi_{12}\\ \xi_{21}&\xi_{22}\end{bmatrix}\in GL_{2}(\mathbb{Q}).$
Let
$\phi(x,y)=\rho_{1}(\xi_{11}x+\xi_{12}y)+\rho_{2}(\xi_{21}x+\xi_{22}y).$ (1)
As a basic example, all non-degenerate rational quadratic forms can be written in the form of (1).
Note that $\xi_{11}\xi_{22}-\xi_{12}\xi_{21}\neq 0$. We say that $\phi(x,y)$ is diagonal if exactly two of the entries $\xi_{11}$, $\xi_{12}$, $\xi_{21}$, $\xi_{22}$ are zero. In such a situation, one has $\phi(x,y)=\tilde{\rho}_{1}(x)+\tilde{\rho}_{2}(y)$ for some polynomials $\tilde{\rho}_{1}(x)$ and $\tilde{\rho}_{2}(y)$ of degree $d$. And we say that $\phi(x,y)$ is non-diagonal if at most one of the entries $\xi_{11}$, $\xi_{12}$, $\xi_{21}$, $\xi_{22}$ is zero.
@@
Let $\phi$ be given as in (1). Let $0<s<1$ and $C_{1},C_{2},c>0$. There exist positive constants $\epsilon=\epsilon(d,s)$, $\epsilon^{\prime}=\epsilon^{\prime}(d,\Xi,s)$ and an integer $N=N(\phi,s,C_{1},C_{2},c)$ such that the following statement holds. Let $(X,Y)$ be a bivariate real random variable whose range is contained in $c,c]^{2}$. Suppose that $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, with $1\ll_{s}s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$. Then, for all $n>N$,
@@
Here, a constant depends on the matrix $\Xi$ means that it depends on the entries of $\Xi$. And a constant depends on $\phi$ means it depends on $\rho_{1}$, $\rho_{2}$ and $\Xi$.
@@
Let $\phi$ be given as in (1). Let $0<s<1$ and $C_{1},C_{2},c>0$. There exist positive constants $\epsilon=\epsilon(s)$, $\epsilon^{\prime}=\epsilon^{\prime}(\phi,s)$ and an integer $N=N(\phi,s,C_{1},C_{2},c)$ such that the following statement holds. Let $X$ and $Y$ be two independent real random variables whose ranges are contained in [c,c][-c,c]. Suppose that $(X,Y)$ is marginally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, with $1\ll_{s}s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$. Then, for all $n>N$,
@@
    The positive $\epsilon$ might fail to exist if one of the conditions of the theorem is not satisfied, including when $s=0$ or 11 (see Example 7.1), the Frostman constant $C$ grows with $n$ (see Example 7.4), or the polynomial is degenerate (see Example 7.3).
@@
    The value of $\epsilon$ is not explicit, and our arguments suggest that they are quantitatively small.
We record several further remarks. Our proof of Theorems 1.21.4 presents a connection between $\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}$ and the entropy of the “distance” function, i.e. $H_{n}\big((X_{1}-Y_{1})^{2}+(X_{2}-Y_{2})^{2}\big)$, where $(X_{1},Y_{1})$, $(X_{2},Y_{2})$ are i.i.d. copies of $(X,Y)$. When $a_{1}=a_{3}=0$ and $a_{2}=1$, i.e. $\phi(x,y)=xy$, Máthé and O’Regan [12] established a different type of connection by replacing the entropy of the “distance” function with that of the “slope” function. By combining this with a generalization of Marstrand’s theorem due to Orponen, Shmerkin, and Wang in [21], they were able to prove Theorem 1.4 for the full range of $s$ for this specific bilinear form function. We do not see how their method could be adapted to the form of $\phi$ with $a_{1},a_{3}\neq 0$. Moreover, in terms of applications, i.e. sum-product type problems over $\mathbb{R}$ or $\mathbb{F}_{p}$, the two functions $xy$ and $a_{1}x^{2}+a_{2}xy+a_{3}y^{2}$ appear to exhibit distinct behaviors. It would be interesting to explore more connections between $\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}$ and the entropy of other functions.
We now discuss applications of the entropy theorems.
For $A\subseteq\mathbb{R}$ and $\delta>0$, by $\mathcal{E}_{\delta}(A)$ we mean the $\delta$-covering number of $A$.
@@
$\rho_{1}(A)+_{G}\rho_{2}(A)=\{\rho_{1}(x)+\rho_{2}(y):\,(x,y)\in G\}.$
@@
Let $d\geq 2$ and $\rho_{1}$, $\rho_{2}$ be real polynomials of degree $d$. Let $0<s<1$ and $C\geq 1$. There exist constants $\epsilon=\epsilon(d,s)>0$ and $\delta_{0}=\delta_{0}(\rho_{1},\rho_{2},s,C)>0$ such that the following holds for all $0<\delta\leq\delta_{0}$. Let $A\subseteq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}[0,1]}$ be any $\delta$-separated set of cardinality $\lceil\delta^{-s}\rceil$. Suppose that for all intervals $I$ of length $l$ at least $\delta$, we have the non-concentration estimate
@@
$\mathcal{E}_{\delta}(A+_{G}A)+\mathcal{E}_{\delta}(\rho_{1}(A)+_{G}\rho_{2}(A))\gg_{\rho_{1},\rho_{2},s}C^{-O_{d,s}(1)}\delta^{-\epsilon}(\#A).$
By Theorem 1.3, extending this result to more general polynomials is possible, but requiring minimum degree constraints (3) and (4).
@@
Let $\phi$ be given as in (1). Let $0<s<1$ and $C\geq 1$. There exists $\epsilon>0$, which depends on $d,s$ (or $d,\Xi,s$) when $\phi(x,y)$ is diagonal (or non-diagonal, respectively), and $\delta_{0}=\delta_{0}(\phi,s,C)>0$ such that the following holds for all $0<\delta\leq\delta_{0}$. Let $A\subseteq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}[0,1]}$ be any $\delta$-separated set of cardinality $\lceil\delta^{-s}\rceil$. Suppose that for all intervals $I$ of length $l$ at least $\delta$, we have the non-concentration estimate
@@
$\{y:\,(x,y)\in G\}\geq\delta^{\epsilon}(\#A)\quad\text{for any}\quad x\in A,$ (3)
$\{x:\,(x,y)\in G\}\geq\delta^{\epsilon}(\#A)\quad\text{for any}\quad y\in A,$ (4)
@@
$\mathcal{E}_{\delta}(A+_{G}A)+\mathcal{E}_{\delta}(\phi_{G}(A,A))\gg_{\phi,s}C^{-O_{d,s}(1)}\delta^{-\epsilon}(\#A).$
In contrast to Theorem 1.6, the next result, an application of Theorem 1.4, holds for every scale $\delta>0$ and is stable under taking subsets: it applies to any $B\subseteq A$ with $B\geq\delta^{\varepsilon}(\#A)$.
@@
Let $\phi$ be given as in (1). Let $0<s<1$ and $C\geq 1$. Then there exists $\epsilon>0$, which depends on $d,s$ (or $d,\Xi,s$) when $\phi(x,y)$ is diagonal (or non-diagonal, respectively), such that the following holds for all $\delta>0$. Let $A\subseteq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}[0,1]}$ be any $\delta$-separated set of cardinality $\lceil\delta^{-s}\rceil$. Suppose that for all intervals $I$ of length $l$ at least $\delta$, we have the non-concentration estimate
@@
$\mathcal{E}_{\delta}(B+B)+\mathcal{E}_{\delta}(\phi(B,B))\gg_{\phi,s}C^{-O_{d,s}(1)}\delta^{-\epsilon}(\#A).$